KyledaGreat:

The half-life of silver-105 is approximately 41.3 days. Step 1 of 3 : Determine a so that A(t)=A0a^t describes the amount of silver-105 left after t days, where A0 is the amount at time t=0. Round to six decimal places. a =

KyledaGreat:

is it A(t) = A0a^t = A(0) = A0a^0 ?

KyledaGreat:

@vocaloid

Vocaloid:

took me a little bit if the half-life is 41.3 days, that means at t = 41.3, A(t) is 1/2 of A0 A(t) = A0*a^t (1/2)A0 = A0*a^41.3 dividing both sides by A0 (1/2) = a^41.3 take the logarithm of both sides log(1/2) = log(a^41.3) using the power of logarithms, you can bring that 41.3 exponent down in front of the log log(1/2) = 41.3log(a) divide both sides by 41.3 log(1/2)/41.3 = log(a) evaluate log(1/2)/41.3 on the left side, then raise both sides to the power of 10 to eliminate the log on the right side and isolate a

KyledaGreat:

it turn out it was −0.00728886 by evaluating

Vocaloid:

good -0.00728886 = log(a) raising both sides to base 10 10^(-0.00728886) = a = 0.983357, rounded to 6 places

KyledaGreat:

0.983357

Vocaloid:

yup

KyledaGreat:

do i put it in with 0. or 983357 since it asked for 6 decimal

Vocaloid:

it's a decimal number so you do have to include the leading decimal 0.983357 would be the proper way to do it

KyledaGreat:

okay i see